OpenStudy (anonymous):

how do i find the slope of a tangent line using the function tanx at the point x=2pi/3

OpenStudy (anonymous):

Derivative as follows: \[\frac{ d }{ dx }\left[ \tan x \right]= \sec^2x=1/\cos^2x\] Now cos(x) = -1/2 at x=2pi/3 and so the slope is 4.